Several blastocoel cavities of early embryos from black mice are injected with the cells you chose....

  1. Several blastocoel cavities of early embryos from black mice are injected with the cells you chose. The blastocyst is then transferred to a pseudopregnant white mouse.  You get the following pups.  Which of the following would you use to continue the experiment?

    A brown and black mouse

    A pink mouse

    A white mouse

    A black mouse

1 points   

QUESTION 3

  1. Knowing that the neomycin resistance gene that you used for the knockout construct is 3 kb in length and knowing that your construct replaces the Frizzled gene with the neomycin gene, which of the mice above are the one you want to study for its knock out phenotype?

    A

    B

    C

    None of the above

0.5 points   

QUESTION 4

  1. A specific gene is 6kb (6000 bases) long. It has 3 introns that are 0.5, 0.5 and 0.5 kb in length, respectively.  The transcription +1 site is at base 200 from the 5' end. The A of the AUG site is 300 bases downstream (or away from the 5' end of the gene) from the +1 site and the TAG stop codon is 4.5 kb farther downstream from the AUG. What is the size of the primary structure of the protein produced from this gene (in amino acids)?

    6000

    4500

    3000

    2000

    999

    299

    199
Answer #1

Answer. 2.data is incomplete because there is no data about pups produced. On the basis of pups we can predict the experimental mouse, if the offspring were of different color it will indicate that mouse have their genetic makeup and they do not rely for coat color on pseudo pregnant mother. The genome will be of their genetic parents.

Answer 3. This data is also incomplete because there is no data for what are A, B and C.

Answer 4. Total length of gene = 6Kb

Gene start from +1 at 200 bases but start codon is at 300bp and end at 4.5kb (which means 1500bp is not used)

So total gene = 6000-(300+1500) = 4200kb.

Also intron does not make protein so total 1500kb

Then gene = 4200-1500 = 2700bp

Also we know that codon is made up of 3 bases or 3 bases code for 1 amino acid..

Then 2700/3 = 900 and also one last stop codon does not code for any gene so 899 amino acid